the gravitational field in a region is given by E=(10N/Kg)(i+2j+3k). find work done by an external agent to slowly shift a particle of mass 3kg from point O(0,0,0) TO A POINT P(3m, 4m, 5m).
answer=-780J

Grade:11

1 Answers

Deepak Kumar
6 years ago
Mass of Particle = 3Kg.
Gravitational Field = E = (10i +20j + 30k)N/Kg.
Initial Position = O(0,0,0)m.
Final Position = P(3,4,5)m.
So Force on Particle = E*Mass of partical = (10i +20j +30k)*3= (30i +60j + 90k)N.
Work done by external Force = – work done by Gravitational force.
                                           = – (30i +60j + 90k)* ((3-0)i +(4-0)j +(5-0)K).
                                           = – (90 + 240 + 450)Joule = – 780 Joule.